### 06 Apr 2026 v1

@@ 4 A path-following smoothing Newton method
This section proposes a path-following smoothing Newton method for symmetric cone programming. The method adopts a two-phase structure. In the first phase, an initial point within a well-defined neighborhood of the central path is efficiently constructed. Starting from this point, the second phase iteratively refines a solution within the neighborhood.
@@ 4.1 Neighborhood of the central path
For practical implementation and theoretical analysis, maintaining iterates within a well-defined neighborhood of the central path is critical. To measure the proximity of a point to the central path and drive the iteration process, we introduce some important functions.
By Theorem 3.2, the function $\eta_{\rho}(\hat{x},s;\mu)$ is strictly convex in $\hat{x}\in\hat{\mathbb{E}}$ and strictly concave in $s\in\mathbb{E}$. Accordingly, we measure its sub-optimality by the primal-dual gap function following (14):
$\theta_{\rho}(\hat{x},s;\mu)$ $=\max_{\tilde{s}}\eta_{\rho}(\hat{x},\tilde{s};\mu)-\min_{\tilde{x}}\eta_{\rho}(\tilde{x},s;\mu)$ (74)
@@
where
$s_{\rho}(\hat{x},\mu)={{\operatorname*{arg\,max}_{\tilde{s}}}\,\eta_{\rho}(\hat{x},\tilde{s};\mu)},\ \hat{x}_{\rho}(s,\mu)={{\operatorname*{arg\,min}_{\tilde{x}}}\,\eta_{\rho}(\tilde{x},s;\mu)}.$
Let $(\hat{x}(\mu),s(\mu))$ be the saddle point of the minimax problem (33), which always exists for any $\mu>0$ and $\rho\geq 1$ by Remark 2. Then for any $(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$, the following inequalities hold:
$\max_{\tilde{s}}\eta_{\rho}(\hat{x},\tilde{s};\mu)\geq\min\limits_{\tilde{x}}\max_{\tilde{s}}\eta_{\rho}(\tilde{x},\tilde{s};\mu)=\eta_{\rho}(\hat{x}(\mu),s(\mu);\mu),$ (75)
$\min_{\tilde{x}}\eta_{\rho}(\tilde{x},s;\mu)\leq\max\limits_{\tilde{s}}\min\limits_{\tilde{x}}\eta_{\rho}(\tilde{x},\tilde{s};\mu)=\eta_{\rho}(\hat{x}(\mu),s(\mu);\mu).$
Let ${\rm val}\,(\mathrm{P}_{\mu})$ denote the optimal value of the primal barrier problem (58). By Theorem 3.3,
@@
This combined with (75) implies that, for any $(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$,
$|\eta_{\rho}(\hat{x},s;\mu)-{\rm val}\,(\mathrm{P}_{\mu})|\leq\theta_{\rho}(\hat{x},s;\mu).$ (76)
Moreover, (75) implies that $\theta_{\rho}(\hat{x},s;\mu)\geq 0$. The equality holds if and only if $(\hat{x},s)$ is the saddle point of problem (33). This gap therefore provides a rigorous certificate of optimality and will be used to measure the proximity to the central path.
In practice, evaluating the exact primal-dual gap $\theta_{\rho}(\hat{x},s;\mu)$ is computationally prohibitive, as it requires solving optimization problems. Recall that $x=\bar{x}+\mathcal{B}\hat{x}$. Let $(\Delta x,\Delta s,\Delta\lambda)$ be the search direction given by (45), and let $\Delta w=(\Delta\hat{x},\Delta s)$ be the Newton direction for the minimax problem (33). Define $w:=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$. For algorithmic purposes, we follow (14) and introduce easily computable merit functions:
$\delta_{\hat{x},\rho}(\hat{x},s;\mu)$ $=\sqrt{\dfrac{1}{\mu}\langle\Delta\hat{x},{D_{\hat{x}\hat{x}}^{2}{\eta_{\rho}(\hat{x},s;\mu)}}\Delta\hat{x}\rangle}$, (77a)
$\delta_{s,\rho}(\hat{x},s;\mu)$ $=\sqrt{-\dfrac{1}{\mu}\langle\Delta s,D_{ss}^{2}\eta_{\rho}(\hat{x},s;\mu)\Delta s\rangle},$ (77b)
$\delta_{\rho}(\hat{x},s;\mu)$ $=\sqrt{(\delta_{\hat{x},\rho}(\hat{x},s;\mu))^{2}+(\delta_{s,\rho}(\hat{x},s;\mu))^{2}}=\|\Delta w\|_{\eta_{\rho},w},$ (77c)
$\xi_{\hat{x},\rho}(\hat{x},s;\mu)$ $=\sqrt{\frac{1}{\mu}\langle\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu),D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(\hat{x},s;\mu)^{-1}\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu)\rangle},$ (77d)
$\xi_{s,\rho}(\hat{x},s;\mu)$ $=\sqrt{-\frac{1}{\mu}\langle\nabla_{s}\eta_{\rho}(\hat{x},s;\mu),D^{2}_{ss}\eta_{\rho}(\hat{x},s;\mu)^{-1}\nabla_{s}\eta_{\rho}(\hat{x},s;\mu)\rangle},$ (77e)
$\xi_{\rho}(\hat{x},s;\mu)$ $=\sqrt{(\xi_{\hat{x},\rho}(\hat{x},s;\mu))^{2}+(\xi_{s,\rho}(\hat{x},s;\mu))^{2}}=\|\nabla_{w}\eta_{\rho}(\hat{x},s;\mu)\|^{*}_{\eta_{\rho},w}.$ (77f)
@@
Note that $\Delta\hat{x}=\mathcal{B}^{*}\Delta x$, $\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu)=\mathcal{B}^{*}\nabla_{x}\eta_{\rho}(\hat{x},s;\mu)$, and
$D_{xx}^{2}\eta_{\rho}(\hat{x},s;\mu)=\mathcal{B}\,D_{\hat{x}\hat{x}}^{2}\eta_{\rho}(\hat{x},s;\mu)\,\mathcal{B}^{*}.$
This implies
$\delta_{x,\rho}(x,s;\mu)$ $=\sqrt{\frac{1}{\mu}\langle\Delta x,{D_{xx}^{2}\eta_{\rho}(\hat{x},s;\mu)}\Delta x\rangle}=\delta_{\hat{x},\rho}(\hat{x},s;\mu),$ (78)
$\xi_{x,\rho}(x,s;\mu)$ $=\sqrt{\frac{1}{\mu}\langle\nabla_{x}\eta_{\rho}(\hat{x},s;\mu)D^{2}_{xx}\eta_{\rho}(\hat{x},s;\mu)^{-1}\nabla_{x}\eta_{\rho}(\hat{x},s;\mu)\rangle}$
$\ =\xi_{\hat{x},\rho}(\hat{x},s;\mu)$
In practical computations, it is unnecessary to explicitly form $\Delta\hat{x}$ and $\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu)$. The quantities $\delta_{x,\rho}(x,s;\mu)$ and $\xi_{x,\rho}(x,s;\mu)$ can be computed directly from $\Delta x$ and $\nabla_{x}\eta_{\rho}(\hat{x},s;\mu)$. For the theoretical analysis, however, we work with $\delta_{\hat{x},\rho}(\hat{x},s;\mu)$ and $\xi_{\hat{x},\rho}(\hat{x},s;\mu)$.
Define the following auxiliary merit functions:
$\tilde{\delta}_{\hat{x},\rho}(\hat{x},s;\mu)$ $=\sqrt{\dfrac{1}{\mu}\langle\widetilde{\Delta\hat{x}},D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(\hat{x},s;\mu)\widetilde{\Delta\hat{x}}\rangle}$, (79a)
$\tilde{\delta}_{s,\rho}(\hat{x},s;\mu)$ $=\sqrt{-\dfrac{1}{\mu}\langle\widetilde{\Delta s},D^{2}_{ss}\eta_{\rho}(\hat{x},s;\mu)\widetilde{\Delta s}\rangle}$, (79b)
$\tilde{\delta_{\rho}}(\hat{x},s;\mu)$ $=\sqrt{(\tilde{\delta}_{\hat{x},\rho}(\hat{x},s;\mu))^{2}+(\tilde{\delta}_{s,\rho}(\hat{x},s;\mu))^{2}}=\|\widetilde{\Delta w}\|_{\eta_{\rho},w},$ (79c)
[…]
If $\xi_{\rho}(\hat{x},s;\mu)=0$, then $(\hat{x},s)$ is the saddle point of the minimax problem (33). This defines a central path that coincides with the one generated by IPMs, as shown in Theorem $3.$ Specifically,
@@
$\Longleftrightarrow$ $\quad\mathcal{A}x=b,\,\mathcal{A}^{*}\lambda+s=c,\,\xi_{\rho}(\hat{x},s;\mu)=0.$
Motivated by this, we define the central-path neighborhood for the SNM based on the reduced SBAL function $\eta_{\rho}$ by
$\mathcal{N}(\kappa,\mu,\rho)=\left\{(x,s,\lambda)\in\mathbb{E}\times\mathbb{E}\times\mathbb{R}^{m}\,|\,\mathcal{A}x=b,\,\mathcal{A}^{*}\lambda+s=c,\,\xi_{\rho}(\hat{x},s;\mu)\leq\kappa\right\},$ (81)
@@
This neighborhood differs from the standard neighborhoods used in classical interior-point Monteiro and Zhang (1998); Nesterov and Todd (1998); Schmieta and Alizadeh (2003) or non-interior Burke and Xu (2000, 1998); Chen and Tseng (2003); Zhao and Li (2003) path-following methods in the literature. It is defined via the merit function induced by the minimax problem, and provides a more faithful characterization of the behavior of the SNM. The proposed method follows the standard paradigm of path-following methods. In the first phase, the iterates are driven into $\mathcal{N}(\kappa,\mu^{(0)},\rho)$. In the second phase, the trajectory is confined to $\mathcal{N}(\kappa,\mu^{(k)},\rho)$, which ensures that all subsequent iterates remain well behaved.
@@ 4.2 Two-phase path-following framework
Before introducing the two-phase framework, we present some necessary notations. Let $v:=(x,s,\lambda)$. Define $\Delta v:=(\Delta x,\Delta s,\Delta\lambda)$ as the concatenation of $\Delta x$, $\Delta s$, and $\Delta\lambda$. Let $K(\theta_{\rho}):=\bigl\{(\hat{x},s):\theta_{\rho}(\hat{x},s;\mu)<+\infty\bigr\}$. For a given initialization point $w^{(0,0)}:=(\hat{x}^{(0,0)},s^{(0,0)})\in K(\theta_{\rho})$, let $\bar{x}\in\mathbb{E}$ with $\mathcal{A}\bar{x}=b$, and $x^{(0,0)}=\bar{x}+\mathcal{B}\hat{x}^{(0,0)}$. For $t\in[0,1]$, define the perturbed function
@@
Since $\eta_{\rho}$ is a nondegenerate $\mu$-self-concordant convex-concave function, $\eta_{t,\rho}$ inherits the same property for any $t\geq 0$.
The first-phase algorithm aims to generate a strictly feasible point within the neighborhood $\mathcal{N}(\kappa,\mu^{(0)},\rho)$. The parameter $t$ is used to scale the merit function so that the initial point $(x^{(0,0)},s^{(0,0)},\lambda^{(0,0)})$ lies in the central-path neighborhood. For this purpose, consider the minimax problem
$\min_{\hat{x}}\max_{s}\left\{\eta_{t,\rho}(w;\mu^{(0)})\right\},$
with first-order optimality conditions
@@
In practical computations, the variable $\hat{x}$ is neither explicitly computed nor formulated. Following Theorem 3.4, we instead solve the system in $(x,s,\lambda)$:
@@
The resulting search direction coincides with that obtained by solving the reduced system (84) under the constraint $\mathcal{A}x=b$. This equivalence is formalized in the following corollary.
@@
Let $(x,s,\lambda)$ satisfy the linear constraint $\mathcal{A}x=b$. For the given point $(x^{(0,0)},s^{(0,0)},\lambda^{(0,0)})$ and any $\mu^{(0)}>0,\rho\geq 1$, suppose that $(\Delta x,\Delta s,\Delta\lambda)$ solves (85). Then the pair $(\Delta\hat{x},\Delta s):=(\mathcal{B}^{*}\Delta x,\Delta s)$ is the unique solution of (84). Conversely, if $(\Delta\hat{x},\Delta s)$ solves (84), define
@@
We start from the initial point $(x^{(0,0)},s^{(0,0)},\lambda^{(0,0)})$ satisfying $\mathcal{A}\bar{x}=b$. Choose $t^{(0)}$ such that
$(1-t^{(0)})\delta_{\rho}(w^{(0,0)};\mu^{(0)})\leq\frac{\kappa}{2}.$ (87)
Repeatedly solving the linear system (85) while updating $t$, we eventually obtain a point $(x^{(0,j)},s^{(0,j)},\lambda^{(0,j)})$ lying in the neighborhood $\mathcal{N}(\kappa,\mu^{(0)},\rho)$. The corresponding algorithmic framework is summarized below.
@@
Step 1: Choose $(\hat{x}^{(0,0)},s^{(0,0)})\in K(\theta_{\rho})$, $\lambda^{(0,0)}\in\mathbb{R}^{m}$, $\rho\geq 1,\mu^{(0)}>0$, and $t^{(0)}\in[0,1]$
     satisfying (87). Set $j:=0$.
Step 2: Compute $\delta_{\rho}^{(0,j)}:=\delta_{\rho}({w}^{(0,j)};\mu^{(0)})$. If $\delta^{(0,j)}_{\rho}\leq\kappa$, compute the Newton
     direction $\Delta v^{(0,j)}$ from (45), set
$v^{(0)}=v^{(0,j)}+\Delta v^{(0,j)},$
     and terminate. Otherwise go to Step 3.
Step 3: Update $t^{(j+1)}=(1-\alpha^{(j)})t^{(j)}$, where
$\alpha^{(j)}=\min\left\{\frac{\kappa}{4t^{(j)}{\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)};\mu^{(0)})^{-1}\nabla_{w}\eta_{\rho}(w^{(0,0)};\mu^{(0)})}\|_{\eta_{\rho},w^{(0,j)}}},1\right\}.$ (88)
Step 4: Compute the search direction $\Delta v^{(j)}$ from (85) with $t=t^{(j+1)}$. Set
$v^{(0,j+1)}=v^{(0,j)}+\Delta v^{(j)}.$ (89)
     Set $j:=j+1$ and return to Step 2.
Throughout Algorithm 1, the iterates $(x^{(0,j)},s^{(0,j)},\lambda^{(0,j)})$ always satisfy the primal constraint by (85). When Algorithm 1 terminates, the second equation of (45) yields
@@
Moreover, Theorem 2.3(iii) gives
$\xi_{\rho}(w^{(0)};\mu^{(0)})\leq\dfrac{\delta^{(0,j)}_{\rho}}{2}\leq\kappa,$ (90)
which implies that $v^{(0)}\in\mathcal{N}(\kappa,\mu^{(0)},\rho)$.
Once a point in $\mathcal{N}(\kappa,\mu^{(0)},\rho)$ has been obtained, we decrease the barrier parameter and, for each new value of the parameter, apply Newton steps to re-enter the corresponding neighborhood. The Newton direction is generated by (45). In what follows, we present the second-phase algorithm.
@@
Step 1: Input $v^{(0,0)}:=v^{(0)}\in\mathcal{N}(\kappa,\mu^{(0)},\rho)$, $\rho\geq 1$, and $\mu^{(0)}>0$. Set $k:=0$ and $j:=0$.
Step 2: Compute the search direction $\Delta v^{(k,j)}$ from (45) and update
$v^{(k,j+1)}=v^{(k,j)}+\Delta v^{(k,j)}.$
Step 3: Compute $\xi^{(k,j+1)}_{\rho}:=\xi_{\rho}(\hat{x}^{(k,j+1)},s^{(k,j+1)};\mu^{(k)})$. If $\xi^{(k,j+1)}_{\rho}>\kappa$, set $j:=j+1$
     and go to Step 2. Otherwise, go to Step 4.
Step 4: Set $v^{(k)}:=v^{(k,j+1)}$ and $\xi^{(k)}_{\rho}:=\xi^{(k,j+1)}_{\rho}$. If $\mu^{(k)}>\varepsilon$, update $\mu^{(k+1)}=\sigma\mu^{(k)}$,
     $v^{(k+1,0)}=v^{(k)}$, set $k:=k+1$, $j:=0$, and return to Step 2. Otherwise, terminate
     and return $v^{(k)}$.
After each reduction of the barrier parameter, the current iterate $v^{(k)}$ is used as the starting point for the next outer iteration and is recentered until it re-enters the new neighborhood $\mathcal{N}(\kappa,\mu^{(k+1)},\rho)$. The well-definedness of this procedure and the bound on the number of inner Newton steps will be established in the next section.
Some remarks on Algorithms 1 and 2 are in order:
@@
    In practical computations, the variable $\hat{x}$ is never explicitly formed. Computations are performed directly with the primal variable $x$ since all the quantities can be evaluated from $(x,s,\lambda)$; see Remark 3 for details.
@@
    The iterate entering Algorithm 2 satisfies the primal and dual constraints. Consequently, by (45), (85), and the iteration scheme, all subsequent iterates preserve feasibility:
    $\mathcal{A}x^{(k,j)}$ $=b,\,$ $\mathcal{A}^{*}\lambda^{(k,j)}+s^{(k,j)}=c,\quad\forall\,j\geq 0,\,k\geq 1.$ (91)
@@
    $\mu^{(k)}\leq\varepsilon\;\text{ and }\;\xi^{(k)}_{\rho}=\|\nabla_{w}\eta_{\rho}(w^{(k)};\mu^{(k)})\|_{\eta_{\rho},w^{(k)},\mu^{(k)}}^{*}\leq\kappa.$ (92)
    It follows from the definition of $\xi^{(k)}_{\rho}$ and Corollary 1 that
    $\|z^{(k)}_{\rho}-x^{(k)}\|$ $\leq\sqrt{\langle z^{(k)}_{\rho}-x^{(k)},\mathcal{H}(z^{(k)}_{\rho}-x^{(k)})\rangle}$ (93)
    $\leq\sqrt{\frac{\mu^{(k)}}{\rho}}\|\nabla_{w}\eta_{\rho}(w^{(k)},\mu^{(k)})\|^{*}_{\eta_{\rho},w^{(k)},\mu^{(k)}}$

### 21 May 2026 v2

@@ 4 A path-following smoothing Newton method
This section proposes a path-following smoothing Newton method for symmetric cone programming. The method consists of two phases. In the first phase, an initial point is constructed in a well-defined neighborhood of the central path. The second phase then uses this point to generate iterates that remain in the neighborhood and terminates once the prescribed accuracy is reached.
@@ 4.1 Neighborhood of the central path
For both practical implementation and theoretical analysis, maintaining iterates within a well-defined neighborhood of the central path is critical. To measure the proximity of a point to the central path and drive the iteration process, we introduce several auxiliary merit functions.
By Theorem 3.2, the function $\eta_{\rho}(w;\mu)$ is strictly convex in $\hat{x}\in\hat{\mathbb{E}}$ and strictly concave in $s\in\mathbb{E}$. Accordingly, we measure its suboptimality by the primal-dual gap function as in (13):
$\theta_{\rho}(w;\mu)$ $=\max_{\tilde{s}}\eta_{\rho}(\hat{x},\tilde{s};\mu)-\min_{\tilde{x}}\eta_{\rho}(\tilde{x},s;\mu)$ (52)
@@
where $s_{\rho}(\hat{x},\mu)={{\operatorname*{arg\,max}_{\tilde{s}}}\,\eta_{\rho}(\hat{x},\tilde{s};\mu)},\,\hat{x}_{\rho}(s,\mu)={{\operatorname*{arg\,min}_{\tilde{x}}}\,\eta_{\rho}(\tilde{x},s;\mu)}.$
Let $(\hat{x}(\mu),s(\mu))$ be the saddle point of the minimax problem (29), which always exists for any $\mu>0$ and $\rho>0$ by Remark 2. Then for any $(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$, the following inequalities hold:
$\max_{\tilde{s}}\eta_{\rho}(\hat{x},\tilde{s};\mu)\geq\eta_{\rho}(\hat{x}(\mu),s(\mu);\mu)\geq\min_{\tilde{x}}\eta_{\rho}(\tilde{x},s;\mu).$ (53)
Let ${\rm val}\,(\mathrm{P}_{\mu})$ denote the optimal value of the barrier problem (47). By Theorem 3.3,
@@
Combining this with (53) yields that, for any $(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$,
$\left|\eta_{\rho}(w;\mu)-{\rm val}\,(\mathrm{P}_{\mu})\right|\leq\theta_{\rho}(w;\mu).$ (54)
Furthermore, (53) implies that $\theta_{\rho}(w;\mu)\geq 0$, and $\theta_{\rho}(w;\mu)=0$ holds if and only if $(\hat{x},s)$ is the saddle point of (29). This gap therefore provides a certificate of optimality and will be used to measure the proximity to the central path.
In practice, evaluating the exact primal-dual gap $\theta_{\rho}(w;\mu)$ is computationally prohibitive, as it requires solving optimization problems. Recall that $x=\bar{x}+\mathcal{B}\hat{x}$. Let $(\Delta x,\Delta s,\Delta\lambda)$ be the search direction given by (40), and let $\Delta w=(\Delta\hat{x},\Delta s)$ be the Newton direction for the minimax problem (29). For algorithmic purposes, we follow (13) and introduce easily computable merit functions:
$\delta_{\hat{x},\rho}(w;\mu)$ $=\sqrt{\dfrac{1}{\mu}\langle\Delta\hat{x},{D_{\hat{x}\hat{x}}^{2}{\eta_{\rho}(w;\mu)}}\Delta\hat{x}\rangle}$, (55a)
$\delta_{s,\rho}(w;\mu)$ $=\sqrt{-\dfrac{1}{\mu}\langle\Delta s,D_{ss}^{2}\eta_{\rho}(w;\mu)\Delta s\rangle},$ (55b)
$\delta_{\rho}(w;\mu)$ $=\sqrt{(\delta_{\hat{x},\rho}(w;\mu))^{2}+(\delta_{s,\rho}(w;\mu))^{2}}=\|\Delta w\|_{\eta_{\rho},w},$ (55c)
$\xi_{\hat{x},\rho}(w;\mu)$ $=\sqrt{\frac{1}{\mu}\langle\nabla_{\hat{x}}\eta_{\rho}(w;\mu),(D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(w;\mu))^{-1}\nabla_{\hat{x}}\eta_{\rho}(w;\mu)\rangle},$ (55d)
$\xi_{s,\rho}(w;\mu)$ $=\sqrt{-\frac{1}{\mu}\langle\nabla_{s}\eta_{\rho}(w;\mu),(D^{2}_{ss}\eta_{\rho}(w;\mu))^{-1}\nabla_{s}\eta_{\rho}(w;\mu)\rangle},$ (55e)
$\xi_{\rho}(w;\mu)$ $=\sqrt{(\xi_{\hat{x},\rho}(w;\mu))^{2}+(\xi_{s,\rho}(w;\mu))^{2}}=\|\nabla_{w}\eta_{\rho}(w;\mu)\|^{*}_{\eta_{\rho},w}.$ (55f)
@@
By Corollary 1, we have
$\delta_{\hat{x},\rho}(w;\mu)=\sqrt{\frac{1}{\mu}\langle\Delta x,{D_{xx}^{2}\eta_{\rho}(w;\mu)}\Delta x\rangle}.$
This implies that it is unnecessary to explicitly form $\Delta\hat{x}$ and $\nabla_{\hat{x}\hat{x}}^{2}\eta_{\rho}(w;\mu)$ in practical computations. The quantity $\delta_{\hat{x},\rho}(w;\mu)$ can be computed directly from $\Delta x$ and $D^{2}_{xx}\eta_{\rho}(w;\mu)$.
If $\xi_{\rho}(w;\mu)=0$, then $(\hat{x},s)$ is the saddle point of the minimax problem (29). This defines a central path that coincides with the one generated by IPMs, as shown in Theorem 3.3. Specifically,
@@
$\Longleftrightarrow$ $\quad\mathcal{A}x=b,\,\mathcal{A}^{*}\lambda+s=c,\,\xi_{\rho}(w;\mu)=0.$
Motivated by this equivalence, we define the central-path neighborhood for the SNM based on the reduced BAL function $\eta_{\rho}$ by
$\mathcal{N}(\kappa,\mu,\rho):=\left\{(x,s,\lambda)\in\mathbb{E}\times\mathbb{E}\times\mathbb{R}^{m}\,|\,\mathcal{A}x=b,\,\mathcal{A}^{*}\lambda+s=c,\,\xi_{\rho}(w;\mu)\leq\kappa\right\},$ (57)
@@
This neighborhood differs from the standard neighborhoods used in classical interior-point path-following methods Monteiro and Zhang (1998); Nesterov and Todd (1998); Schmieta and Alizadeh (2003) or in non-interior path-following methods Burke and Xu (2000, 1998); Chen and Tseng (2003); Zhao and Li (2003). It is defined via the merit function induced by the minimax problem, and is tailored to the structure of the SNM. The proposed method follows the standard paradigm of path-following methods. In the first phase, the iterates are driven into $\mathcal{N}(\kappa,\mu^{(0)},\rho)$. In the second phase, the iterates are maintained in $\mathcal{N}(\kappa,\mu^{(k)},\rho)$ while the smoothing parameter is updated.
@@ 4.2 Two-phase path-following framework
Before introducing the two-phase framework, we present some additional notation. Let
$v:=(x,s,\lambda),\quad\Delta v:=(\Delta x,\Delta s,\Delta\lambda).$
Define $K(\theta_{\rho}):=\bigl\{w\mid\theta_{\rho}(w;\mu)<+\infty\bigr\}$. For a given initial point $w^{(0,0)}:=(\hat{x}^{(0,0)},s^{(0,0)})\in K(\theta_{\rho})$, let $\bar{x}\in\mathbb{E}$ satisfy $\mathcal{A}\bar{x}=b$, and set
$x^{(0,0)}=\bar{x}+\mathcal{B}\hat{x}^{(0,0)}.$
For $t\in(0,1]$, define the perturbed function
@@
The purpose of this perturbation is to construct a continuation path from the initial point to a point satisfying the neighborhood condition for the fixed parameter $\mu^{(0)}$. Since $\eta_{\rho}$ is a nondegenerate $\mu$-self-concordant convex-concave function, $\eta_{t,\rho}$ inherits the same property for any $t\geq 0$.
The first phase of PFSNM aims to generate a feasible point in the neighborhood $\mathcal{N}(\kappa,\mu^{(0)},\rho)$. The parameter $t$ scales the perturbation term so that the initial point $(x^{(0,0)},s^{(0,0)},\lambda^{(0,0)})$ lies in the central-path neighborhood of the perturbed problem. For this purpose, consider the minimax problem
$\min_{\hat{x}\in\hat{\mathbb{E}}}\max_{s\in\mathbb{E}}\left\{\eta_{t,\rho}(w;\mu^{(0)})\right\}$
with the first-order optimality conditions
@@
In practical computations, the variable $\hat{x}$ is not formed explicitly. Following Theorem 3.4, we instead solve the system in $(x,s,\lambda)$:
@@
The resulting search direction coincides with the one obtained by solving (60) under the constraint $\mathcal{A}x=b$. This equivalence is formalized in the following corollary.
@@
Let $(x,s,\lambda)$ satisfy the affine constraint $\mathcal{A}x=b$. For the given point $(x^{(0,0)},s^{(0,0)},\lambda^{(0,0)})$ and any $\mu^{(0)}>0,\,\rho>0$, suppose that $(\Delta x,\Delta s,\Delta\lambda)$ solves (61). Then the pair $(\Delta\hat{x},\Delta s):=(\mathcal{B}^{*}\Delta x,\Delta s)$ is the unique solution of (60). Conversely, if $(\Delta\hat{x},\Delta s)$ solves (60), define
@@
We are now ready to state the first phase of the proposed method. Starting from an arbitrary point $w^{(0,0)}\in K(\theta_{\rho})$, the purpose of this phase is to follow the perturbed minimax problems defined by $\eta_{t,\rho}$ and to drive the iterate into the neighborhood $\mathcal{N}(\kappa,\mu^{(0)},\rho)$ for the fixed smoothing parameter $\mu^{(0)}$. The parameter $t$ plays the role of a homotopy parameter: for a suitable $t^{(0)}\in(0,1]$ such that
$(1-t^{(0)})\delta_{\rho}(w^{(0,0)};\mu^{(0)})\leq\frac{\kappa}{2},$ (63)
the initial point is close to the saddle point of the perturbed problem, while decreasing $t$ gradually removes the perturbation and recovers the original reduced minimax problem.
@@
[…]
2:for $j=0,1,2,\ldots$ do
3:  if $\delta_{\rho}(w^{(0,j)};\mu^{(0)})\leq\kappa$ then
4:   Compute the Newton direction $\Delta v^{(0,j)}$ from (40).
5:   return $v^{(0)}=v^{(0,j)}+\Delta v^{(0,j)}.$
6:  else
7:   Update
$\quad\alpha^{(j)}=\min\left\{\frac{\kappa}{4t^{(j)}\left\|\bigl(D^{2}_{ww}\eta_{\rho}(w^{(0,j)};\mu^{(0)})\bigr)^{-1}\nabla_{w}\eta_{\rho}(w^{(0,0)};\mu^{(0)})\right\|_{\eta_{\rho},w^{(0,j)}}},1\right\}.$ (64)
8:   Set $t^{(j+1)}=(1-\alpha^{(j)})t^{(j)}.$
9:   Compute the search direction $\Delta v^{(0,j)}$ from (61) with $t=t^{(j+1)}$ and set
$v^{(0,j+1)}=v^{(0,j)}+\Delta v^{(0,j)}.$
10:  end if
11:end for
Throughout Algorithm 1, the iterates $(x^{(0,j)},s^{(0,j)},\lambda^{(0,j)})$ always satisfy the primal constraint, as follows from (61). When Algorithm 1 terminates, the second equation of (40) yields
@@
Furthermore, since $\delta_{\rho}(w^{(0,j)};\mu^{(0)})\leq\kappa=0.1<1$, Theorem 2.3(iii) yields
$\xi_{\rho}(w^{(0)};\mu^{(0)})\leq\left(\frac{\delta_{\rho}(w^{(0,j)};\mu^{(0)})}{1-\delta_{\rho}(w^{(0,j)};\mu^{(0)})}\right)^{2}\leq\frac{1}{2}\delta_{\rho}(w^{(0,j)};\mu^{(0)})\leq\kappa.$
Hence $v^{(0)}\in\mathcal{N}(\kappa,\mu^{(0)},\rho)$.
Once a point in $\mathcal{N}(\kappa,\mu^{(0)},\rho)$ has been obtained, the second phase decreases the smoothing parameter and recenters the iterate for each new value of $\mu$. The analysis in Section 5 will show that, after each update of $\mu$, only one Newton step is sufficient to return to the neighborhood.
@@
[…]
2:for $k=0,1,2,\ldots$ do
3:  if $\mu^{(k)}\leq\varepsilon$ then
4:   return $v^{(k)}$.
5:  else
6:   Update $\mu^{(k+1)}=\sigma\mu^{(k)}$.
7:   Compute the search direction $\Delta v^{(k)}$ from (40) with $\mu=\mu^{(k+1)}$, and set
$v^{(k+1)}=v^{(k)}+\Delta v^{(k)}.$
8:  end if
9:end for
After each reduction of the smoothing parameter, a single full Newton step is sufficient to recenter the iterate; more precisely, the resulting point $v^{(k+1)}$ belongs to the new neighborhood $\mathcal{N}(\kappa,\mu^{(k+1)},\rho)$. This fact is established in Theorem 5.5. Therefore, Algorithm 2 retains the standard path-following structure.
Several remarks on Algorithms 1 and 2 are in order:
@@
    In practical computations, the variable $\hat{x}$ is never formed explicitly. Computations are performed directly with the primal variable $x$ since $\delta_{\rho}(w^{(0,j)};\mu^{(0)})$ can be evaluated from $(x,s,\lambda)$; see Remark 3 for details.
@@
    There is no need to form the Newton systems of $\eta_{\rho}$ or $\eta_{t,\rho}$ with respect to $(\hat{x},s)$ explicitly. By Theorem 3.4 and Corollary 2, updating $(x,s,\lambda)$ is in fact equivalent to updating $(\hat{x},s)$.
@@
    The iterate entering Algorithm 2 satisfies the primal and dual constraints. Consequently, by (40) and the iteration scheme, all subsequent iterates remain feasible:
    $\mathcal{A}x^{(k)}$ $=b,\,$ $\mathcal{A}^{*}\lambda^{(k)}+s^{(k)}=c,\quad\forall\,k\geq 0.$ (65)
  4. (iv)
@@
    $\mu^{(k)}\leq\varepsilon\;\text{ and }\;\xi_{\rho}(w^{(k)};\mu^{(k)})=\|\nabla_{w}\eta_{\rho}(w^{(k)};\mu^{(k)})\|_{\eta_{\rho},w^{(k)},\mu^{(k)}}^{*}\leq\kappa.$
    It follows from the definition of $\xi_{\rho}$ and Corollary 1 that
    $\|z^{(k)}_{\rho}-x^{(k)}\|$ $\leq\sqrt{\langle z^{(k)}_{\rho}-x^{(k)},\mathcal{H}(z^{(k)}_{\rho}-x^{(k)})\rangle}$
    $\leq\sqrt{\frac{\mu^{(k)}}{\rho}}\|\nabla_{w}\eta_{\rho}(w^{(k)};\mu^{(k)})\|^{*}_{\eta_{\rho},w^{(k)},\mu^{(k)}}$